### v1

@@ 2.1.4 Dataset splits
Because each potato tuber was photographed 20 to 30 times, the total number of RGB-D images for training the encoder was 6794. The number of validation images was 1439 and the number of independent test images was 1425.
@@ 2.2.1 Encoder and decoder architecture
Figure 6: Schematic representation of our encoder-decoder network for 3D shape completion of potato tubers. The input data for our network was a four channel RGB-D image that was clipped and masked to the potato region and padded to a fixed dimension of 304 pixels. The input image was fed into the convolutional encoder which compressed the data into a latent vector, whose dimension was configurable. Then, the latent vector was processed by the DeepSDF decoder to reconstruct the complete 3D point cloud.
Our 3D shape completion network was based on the CoRe network presented in Magistri et al. (2022). Given the many updates made to CoRe during our research, our 3D shape completion network was renamed to CoRe++. Figure 6 gives a schematic overview of CoRe++. CoRe++ consisted of a convolutional encoder and a decoder. The encoder was a small neural network consisting of seven convolutional layers, each followed by a Leaky-ReLU activation function and a Max-Pooling layer, see the details in Table 1. The position of the pooling layer was one of the changes of CoRe++ compared to CoRe (in CoRe, the pooling layer was before the activation). We hypothesized that this positional change of the pooling layer could help prevent the loss of important features. After the last convolution block, there was a flatten layer after which the flattened output tensor was fed into a fully connected layer. This layer outputted the latent vector, whose dimension was made configurable for the purpose of our study. Conceptually, the latent vector is a compressed representation of the 3D shape of the potato tuber.
@@ 16 Fully Connected - - - - 131,104
Figure 7: Visualization of the signed distances as target values for DeepSDF to learn the 3D shape of potato tubers. The red colored points are negative distance values representing the 3D points inside the tuber’s surface. The green colored points are positive distance values representing the 3D points outside the tuber’s surface. The blue colored points have a signed distance of zero and represent the points on the target 3D shape for DeepSDF to learn.
@@ 2.2.2 Data input, preprocessing, augmentation, and postprocessing
The input for the CoRe++ network was a four-channel RGB-D image that was masked to the potato region and clipped to a fixed-sized image of 304 x 304 pixels (Figure 6). The image dimension was chosen such that the largest tuber in our dataset would completely fit in the image. The input boxes and masks were obtained after manually annotating the potato tuber regions with the LabelMe annotation software. Note that this clipping and masking can also be achieved with an object detection or instance segmentation model.
Two new data preprocessing techniques were added to CoRe++. The first was a depth pixel filtering algorithm, which was extracted from Blok et al. (2021), that removed all depth pixels that were further away from the depth values of the majority of the pixels in the masked depth image. This majority of pixels was assumed to represent the potato region, and helped to remove depth outliers in the depth image. The second preprocessing technique was a normalization of the remaining depth pixels between the minimum (230 mm) and maximum distance value (350 mm) between the RGB-D camera and the potato tubers on the conveyor belt. This normalization helped to increase contrast in the depth image, which potentially leads to better model convergence.
In addition to these new data preprocessing techniques, we also added geometric and color data augmentations to the encoder training procedure. The geometric augmentations involved a random image rotation to a maximum of 45 degrees, a random horizontal flipping of the image and a random vertical flipping of the image. The geometric data augmentations were applied to both the RGB image and the depth image. The color augmentations were only applied to the RGB image, and involved a random change of the brightness, saturation and hue of the image. The parameters (0.5 for brightness, 0.5 for saturation and -0.1 to 0.1 for hue) were chosen such that the augmented image looked different but visually realistic compared to the original image.
@@
As Figure 6 indicates, the output of the CoRe++ network is a completed 3D point cloud. Since it is not possible to calculate volume directly from a 3D point cloud, we added a 3D data postprocessing procedure that converted the 3D point cloud into a watertight 3D mesh from which the volume can be estimated. Our 3D mesh generation procedure consisted of selecting the predicted signed distance values less than or equal to 0.0. These values represented the 3D points on the surface and inside the potato tuber, refer to the blue- and red-colored points in Figure 7. From the selected points, a 3D convex hull shape was extracted using Open3D’s GPU-accelerated hull function. The convex hull was then smoothed with a triangle-based algorithm (Loop, 1987) that divided each triangle of the hull into four smaller triangles. What followed were noise suppression procedures and an iterative voxel-based downsampling of the generated hull in case it was not watertight. The latter procedure guaranteed that each produced 3D mesh was watertight so that the volume could be calculated.
@@ 2.2.3 Training procedure
At the start of the DeepSDF training, a randomly initialized latent vector was assigned to each data point. These latent vectors were then optimized along with the weights of the decoder using standard backpropagation. In our research, the DeepSDF decoder was trained for 1001 epochs with a step-based learning rate scheduler that started at a learning rate of 5$\cdot$10-4. For every 300 epochs, the learning rate was halved. The decoder was trained with the Adaptive Moment Estimation (Adam) optimizer. We used the L1 loss function during training, because this loss function serves the purpose of minimizing the sum of all distance differences between the target 3D shape and the predicted 3D shape. For every 10 epochs, a weight file and a latent vector code were automatically saved. After training, the weight file with the best 3D reconstruction on the shapes of the validation set was selected. This selected weight file and the corresponding latent vector code were used as targets for training the encoder. The metric for determining the best 3D reconstruction of the DeepSDF decoder was the Chamfer distance. The Chamfer distance ($d_{CD}$) represents the sum of the average closest distance from points in the ground truth 3D shape ($\mathcal{G}$) to the points in the reconstructed 3D shape ($\mathcal{S}$), and vice versa (Equation 1). The lower the Chamfer distance, the better the 3D reconstruction.
@@
For every epoch the encoder was inferred on the validation set with the decoder activated. The encoder weights with the lowest root mean squared error (RMSE, Equation 4) on the volume were stored for final evaluation on the independent test set. We chose the RMSE metric, because it was the most common metric in the scientific literature for evaluating the volumetric estimate of potato tubers. The ground truth volume ($V$, Equation 4) was extracted from the reconstructed 3D mesh (Section 2.1.3). This volume estimate was more accurate than the volume estimate of the ground truth method with water displacement, as the latter had an accuracy of 10 ml at best and was more prone to human reading and writing errors.
@@
2.3 Evaluation
The Chamfer distance (Equation 1, Section 2.2.3) was used as the performance metric for evaluating the 3D point cloud completion. Besides the Chamfer distance, we also calculated the precision, recall and the F-score. The precision ($p$) was the percentage of reconstructed points within a certain distance ($d$) to the ground truth point cloud (Equation 5). As such, it represented the accuracy of the 3D reconstruction. The recall ($r$) was the percentage of ground truth points within a certain distance ($d$) to the reconstructed point cloud (Equation 6). The recall represented the completeness of the 3D reconstruction. For our evaluation, we used a distance ($d$) threshold of 5.0 mm, as this value allowed to compare our results one-on-one with those of Magistri et al. (2022). After calculating the precision and recall, the F-score was obtained (Equation 7). The F-score was the harmonic mean between the precision and the recall, and it represented the percentage of 3D points that were correctly reconstructed.
[…]
For evaluating the volumetric estimate we used the RMSE metric, as presented in Equation 4. To evaluate the processing speed, we calculated the average 3D shape completion time on the test images. This analysis was performed with a Lenovo Legion Pro 7 16IRX8H laptop with a NVIDIA GeForce RTX 4090 Laptop GPU.
@@ 2.4.1 The effect of the latent size on the 3D shape completion result
For 3D shape completion, it is important that the encoder-compressed latent vector contains high-level distinguishable 3D features that generalize well on new and untrained shapes. Because the latent vector plays such an important role in an encoder-decoder network like ours, we set up an experiment that tested what was the optimal size for the latent vector for completing the 3D shape of potato tubers. We investigated six sizes: 8, 16, 32, 64, 128, and 256. Latent size 32 was used in the original study by Magistri et al. (2022). Our hypothesis was that the smaller sizes, such as 8 and 16, had insufficient capacity to compress the 3D shape, leading to suboptimal generalization. Also, we hypothesized that the larger sizes, such as 128 and 256, were probably too large resulting in insufficient compression to generalizable 3D shape features. Hence, we hypothesized that the medium sizes, 32 and 64, were probably more optimal for the compression and thus generalization of 3D shapes.
@@
To better understand the optimal size of the latent vector, we conducted an additional experiment with our CoRe++ network. This experiment consisted of interpolating the values of the latent vector of the smallest potato tuber and the largest in the test set. By interpolating the values between these two latent vectors, we were able to better understand the diversity of the generated 3D shapes and sizes in the corresponding latent space of the six tested latent sizes. Ideally, the latent space is constructed so that after interpolation, the generated 3D shapes are realistic in shape and have a sequential build-up in size, as the interpolation was performed between the smallest and largest tuber.
@@
The experiment was tested with 3 networks: the DeepSDF decoder-only architecture without the encoder, the original CoRe network by Magistri et al. (2022), and our CoRe++ network. The DeepSDF decoder-only architecture was tested to better understand where the largest effect of latent size was: in the encoder or in the decoder. Testing the original CoRe network by Magistri et al. (2022) provided insight into a potential trend commonly shared between CoRe and CoRe++. It also gave a more detailed overview on the potential improvement of CoRe++ over CoRe.
@@
The performance parameters were the ones listed in Section 2.3. Our evaluation was based on two requirements: accuracy and analysis speed. The accuracy requirement was met if the RMSE on the volume was lower than that of a standard linear regression model. Input to that linear regression model were the length, width and depth estimates of the partially completed point cloud (Figure 8). This point cloud was obtained by converting the filtered RGB-D image (Figure 6) into 3D points in Open3D software using the intrinsic camera parameters of the RGB-D camera. From the generated point cloud, the oriented 3D bounding box was obtained from which the length, width and depth dimensions were extracted as input parameters for training the linear regression model. The linear regression model was trained on the same train and validation set and tested on the same independent test set (Section 2.1.4). The trained linear regression model resulted a RMSE of 31.1 ml on the volumetric estimate of the test images. As such, if the 3D shape completion method had a RMSE lower than 31.1 ml, then the accuracy requirement was met.
@@
Figure 8: To determine whether the accuracy requirement was met, we trained a linear regression model on the length, width and depth estimates of the partially completed point cloud. The dimension estimates were obtained from the oriented 3D bounding box which is shown in red.
@@
The analysis speed requirement was met if the 3D shape completion was finished in less than 16 milliseconds (ms) per potato tuber. This analysis time was calculated based on the highest number of potato tubers present in a single image in our dataset: 59. The average throughput time of the potato tubers on the harvester was 0.95 seconds, and this value was derived by dividing the average number of 28.5 frames per tuber (9658/339, Section 2.1.4) by the camera acquisition rate of 30 frames per second. With the maximum number of tubers in a single image and the average throughput time, we estimated that the 3D shape completion method should analyze up to 62 potato tubers per second (59/0.95). This equaled to 16 ms per tuber (1000/62).
@@
The goal of this experiment was to gather insights into the performance of our CoRe++ shape completion method when applied in the field. The first practical analysis was on the effect of the tuber size on the 3D shape completion result. This analysis is important because it provides information about the degree of generalizability of our method in a particular field. Under typical field conditions, a wide variety of tuber sizes move over the harvester’s conveyor belt, which means that our 3D completion method should be able to generate accurate volumetric estimates regardless of the tuber size. We decided to perform the analysis on four different size classes, which were extracted from our dataset distribution (Figure 5): small tubers with a volume between 0 and 100 ml, small to medium tubers with a volume between 100 and 150 ml, medium to large tubers with a volume between 150 and 200 ml, and large tubers with a volume between 200 and 500 ml. The ranges were chosen so that a relatively balanced number of tubers ended up in the four classes.
@@
The second practical analysis was on the effect of the potato cultivar on the 3D shape completion result. This analysis provided insights into the degree of generalizability of the 3D completion method at the farm level, as the majority of farmers grow multiple cultivars in a season. We wanted to test if our method can provide accurate volumetric estimates regardless of the potato cultivar. We conducted the experiment with three cultivars (Corolle, Kitahime and Sayaka) that differed in shape and size. Corolle was the cultivar with the most elongated and smallest tubers. Kitahime was the cultivar with the most spherical and medium-sized tubers. Sayaka was the cultivar with the largest tubers and the greatest diversity in shape. This may be due to the fact that Sayaka was the cultivar with the most samples in our dataset. As shape metrics, we calculated the elongation factor and the concavity factor. The elongation factor was obtained by dividing the longest dimension of the 3D bounding box of the ground truth mesh by the shortest dimension. Spherical shapes have an elongation factor close to 1 and elongated shapes have an elongation factor closer to 2. The concavity factor was obtained by calculating the Chamfer distance between the original ground truth 3D mesh and the convex hull of that mesh. Higher Chamfer distances mean that there are more valleys or concave regions on the surface of the potato tuber.
@@
The final practical analysis was on the effect of the image analysis region on the 3D shape completion result. This analysis provided insight into the RGB-D image region where ideally the 3D shape completion should be performed. This analysis was necessary because it is computationally intensive to analyze all RGB-D frames for each tuber. We intended to identify an area in the RGB-D image where the RMSEs for the volumetric estimate are lowest on average. That particular area will be most promising for performing the high-throughput 3D shape completion. In our analysis, the RMSEs were summarized for thirteen horizontal regions in the RGB-D image (Table 2).
Table 2: Summary of the investigated horizontal pixel regions for analyzing the effect of the image analysis region on the 3D shape completion result.
Region Horizontal region

### v3

@@ 2.1.4 Dataset splits and volume
Because each potato tuber was photographed 20 to 30 times on the conveyor belt of the harvester, the total number of RGB-D images for training the encoder was 6794. The number of validation images was 1439 and the number of independent test images was 1425.
The average data volume per potato was 58.5 MB. This consisted of an average of 55.1 MB for the RGB-D image frames and an average of 3.5 MB for the 3D mesh.
@@ 2.2.1 Encoder and decoder architecture
Figure 6: Schematic representation of our encoder-decoder network for 3D shape completion of potato tubers. The encoder compresses RGB-D images into a latent vector, which is processed by the DeepSDF decoder to reconstruct the complete 3D point cloud.
Our 3D shape completion network was based on the CoRe network presented in Magistri et al. (2022). Given the many updates made to CoRe during our research, our 3D shape completion network was renamed to CoRe++. Figure 6 gives a schematic overview of CoRe++. CoRe++ consisted of a convolutional encoder and a decoder. This encoder-decoder network was chosen over alternatives like 3D convolutions or voxel-based networks due to its small network architecture, thereby enabling high-throughput processing. The encoder was a small neural network consisting of seven convolutional layers, each followed by a Leaky-ReLU activation function and a Max-Pooling layer, see the details in Table 1. The position of the pooling layer was one of the changes of CoRe++ compared to CoRe (in CoRe, the pooling layer was before the activation). We hypothesized that this positional change of the pooling layer could help prevent the loss of important image features. After the last convolution block, there was a flatten layer that flattened the output tensor so that it could be processed by the fully connected layer. The fully connected layer outputted the latent vector, which was a compressed representation of the 3D shape of the potato tuber. The size of the latent vector, which could be configured in our software, determined the network’s ability to optimize the dataset. Intuitively, the larger the size, the higher the number of parameters in the network, meaning that it required larger datasets to properly train.
@@ 16 Fully Connected - - - - 131,104
Figure 7: Visualization of the signed distances as target values for DeepSDF to learn the 3D shape of potato tubers: red points are negative (inside surface), green points are positive (outside surface), and blue points are zero (on the surface).
@@ 2.2.2 Data input, preprocessing, augmentation, and postprocessing
The input for the CoRe++ network was a four-channel RGB-D image that was masked to the potato region and clipped to a fixed-sized image of 304 x 304 pixels (Figure 6). This image dimension was chosen such that the largest tuber in our dataset would completely fit in the image. The input boxes and masks were obtained after manually annotating the potato tuber regions with the LabelMe annotation software. Note that this region clipping and masking can also be achieved with an object detection or instance segmentation model.
Two new data preprocessing techniques were added to CoRe++. The first was a depth pixel filtering algorithm, which was extracted from Blok et al. (2021), that removed all depth pixels that were further away from the depth values of the majority of the pixels in the masked depth image. This majority of pixels was assumed to represent the potato region, and helped to remove depth outliers in the depth image. The second preprocessing technique was a normalization of the remaining depth pixels between the minimum (230 mm) and maximum distance value (350 mm) between the RGB-D camera and the potato tubers on the conveyor belt. This normalization helped to increase contrast in the depth image (Figure 6), which potentially leads to better model convergence.
In addition to these new data preprocessing techniques, we also added geometric and color data augmentations to the encoder training procedure in an attempt to improve the network’s generalization performance. Both geometric and color augmentations were parameterized so that the augmented images looked different but visually realistic compared to the original image. The geometric augmentations involved a random image rotation to a maximum of 45 degrees, a random horizontal flipping of the image and a random vertical flipping of the image. These augmentations mimicked different orientations of the tubers on the conveyor belt. The geometric data augmentations were applied to both the RGB image and the depth image. The color augmentations were only applied to the RGB image, and involved a random change of the brightness, saturation and hue of the image, by using the following parameters: between 0.0 and 0.5 for brightness and saturation, and between -0.1 to 0.1 for hue. These color augmentations mimicked different tuber colors and lighting conditions, helping the network to better generalize to new potato cultivars and different lighting settings.
@@
As Figure 6 indicates, the output of the CoRe++ network is a completed 3D point cloud. Since it is not possible to calculate volume directly from a 3D point cloud, we added a 3D data postprocessing procedure that converted the 3D point cloud into a fully enclosed, watertight 3D mesh from which the volume could be estimated. Our 3D mesh generation consisted of selecting the predicted signed distance values less than or equal to 0.0. These values represented the 3D points on the surface and inside the potato tuber, refer to the blue- and red-colored points in Figure 7. From the selected points, a 3D convex hull shape was extracted using Open3D’s GPU-accelerated hull function. The convex hull was then smoothed with a triangle-based algorithm (Loop, 1987) that divided each triangle of the hull into four smaller triangles. The value of four turned out to be optimal for smoothing the 3D shape in a high-throughput fashion. For applications that are less time-critical, a value higher than four can be chosen to make the 3D shape completion even more accurate. What followed were noise suppression procedures and an iterative voxel-based downsampling of the generated hull in case it was not watertight. The latter procedure guaranteed that each produced 3D mesh was watertight so that the volume could be calculated.
@@ 2.2.3 Training procedure
At the start of the DeepSDF training, a randomly initialized latent vector was assigned to each data point. These latent vectors were then optimized along with the weights of the decoder using standard backpropagation. In our research, the DeepSDF decoder was trained for 1001 epochs with a step-based learning rate scheduler that started at a learning rate of 5$\cdot$10-4 (this value was adopted from the original DeepSDF paper). For every 300 epochs, the learning rate was halved. The decoder was trained with the Adaptive Moment Estimation (Adam) optimizer. We used the L1 loss function during training, because this loss function serves the purpose of minimizing the sum of all distance differences between the target 3D shape and the predicted 3D shape. For every 10 epochs, a weight file and a latent vector code were automatically saved. The snapshot value of 10 epochs was significantly lower than the original DeepSDF implementation (1000) and Magistri et al. (2022) (500), and it allowed us to frequently identify trends of network overfitting, thereby allowing us to make a better selection of the optimal weights. The metric for selecting the optimal weights was the Chamfer distance ($d_{CD}$), which is the sum of the average closest distance from points in the ground truth 3D shape ($\mathcal{G}$) to the points in the reconstructed 3D shape ($\mathcal{S}$), and vice versa (Equation 1). The lower the Chamfer distance, the better the 3D reconstruction. The unit of the Chamfer distance was millimeter (mm). The weight file and corresponding latent vector code with the lowest Chamfer distance on the validation set were used as targets for training the encoder.
@@
For every epoch, the encoder was inferred on the validation set with the decoder activated. The encoder weights with the lowest root mean squared error (RMSE, Equation 4) on the volume were stored for final evaluation on the independent test set. We chose the RMSE metric, because it was the most common metric in the scientific literature for evaluating the volumetric estimate of potato tubers. The ground truth volume ($V$) was calculated from the reconstructed 3D mesh (Section 2.1.3), using the Tetrahedron method (Equation 5). The same method was used to calculate the volume from CoRe++’s 3D reconstruction ($\hat{V}$).
@@
$V=\sum_{i=1}^{T}\frac{1}{6}\left(\mathbf{v}_{1}\cdot(\mathbf{v}_{2}\times\mathbf{v}_{3})\right)$ (5)
where $\mathbf{v}_{1}$, $\mathbf{v}_{2}$, $\mathbf{v}_{3}$ are the three vertices representing a triangle in the mesh. $T$ is the total number of triangles in the mesh.
@@ 2.3.1 The effect of the latent size on the 3D shape completion result
For 3D shape completion, it is important that the encoder-compressed latent vector contains high-level distinguishable 3D features that generalize well on new and untrained shapes. Because the latent vector plays such an important role in an encoder-decoder network like ours, we set up an experiment that tested what was the optimal latent vector size for completing the 3D shape of potato tubers. We investigated six sizes: 8, 16, 32, 64, 128, and 256. Latent size 32 was used in the original study by Magistri et al. (2022). Our hypothesis was that the smaller sizes, such as 8 and 16, compressed the data too much, possibly losing critical details, leading to suboptimal generalization. The larger sizes, such as 128 and 256, potentially captured more noise, leading to overfitting and suboptimal shape generalization. The medium sizes, 32 and 64, probably struck a better balance between compression and generalizability because they allowed the model to efficiently capture key features.
@@
To better understand the optimal size of the latent vector, we conducted an additional experiment with our CoRe++ network. For each of the six tested latent sizes, we interpolated the values of the latent vector of the smallest potato tuber and the largest in the test set. Interpolating between these two latent vectors can reveal how well the network has learned to represent size variations. A smooth transition in the generated 3D shapes and sizes can indicate that the latent space is well-structured, meaning that the network can effectively generate realistic 3D shapes for unseen inputs, i.e., generalize better (Mi et al., 2021).
@@
The experiment was tested with 3 networks: the DeepSDF decoder architecture without the encoder, the original CoRe network of Magistri et al. (2022), and our CoRe++ network. The DeepSDF decoder-only architecture was tested to better understand where the largest effect of latent size was: in the encoder or in the decoder. Testing the original CoRe network of Magistri et al. (2022) provided insight into a potential trend commonly shared between CoRe and CoRe++. It also gave a more detailed overview on the potential improvement of CoRe++ over CoRe.
@@
Our evaluation was based on two requirements: accuracy and analysis speed. The accuracy requirement was met if the RMSE on the volume (Equation 4) was lower than that of a standard linear regression model. Input to that linear regression model were the length, width and depth estimates of the partially completed point cloud (Figure 8). This point cloud was obtained by converting the filtered RGB-D image (Figure 6) into 3D points using the intrinsic camera parameters of the RGB-D camera in Open3D software. From the generated point cloud, the oriented 3D bounding box was obtained from which the length, width and depth dimensions were extracted (Figure 8) as input parameters for training the linear regression model. The linear regression model was trained on the same train and validation set and tested on the same independent test set (Section 2.1.4). Our trained linear regression model resulted a RMSE of 31.1 ml on the volumetric estimate of the 1425 test images. As such, if the 3D shape completion method had a RMSE lower than 31.1 ml, then the accuracy requirement was met.
@@
Figure 8: To determine whether the accuracy requirement was met, we trained a linear regression model on the length, width and depth estimates of the oriented 3D bounding box generated from the partially completed point cloud.
@@
The analysis speed requirement was met if the 3D shape completion was finished in less than 16 milliseconds (ms) per potato tuber. This maximum analysis time was calculated based on the highest number of potato tubers present in a single image in our dataset: 59. The average throughput time of the potato tubers on the harvester was 0.95 seconds, and this value was derived by dividing the average number of 28.5 frames per tuber (9658/339, Section 2.1.4) by the camera acquisition rate of 30 frames per second. With the maximum number of tubers in a single image and the average throughput time, we estimated that the 3D shape completion method should analyze up to 62 potato tubers per second (59/0.95). This equaled to 16 ms per tuber (1000/62).
@@
The goal of this experiment was to gather insights into the performance of our CoRe++ shape completion method when applied in the field. The analysis was conducted with the best performing latent size found in the previous experiment. Our first practical analysis was on the effect of the tuber size on the 3D shape completion result. This analysis is important because it provides information about the degree of generalizability of our method in a particular field. Under typical field conditions, a wide variety of tuber sizes move over the harvester’s conveyor belt, and this size variability can influence yield estimation. Our method must handle this variability effectively, ensuring accurate volumetric estimates regardless of tuber size. Such flexibility is crucial for making reliable yield predictions in real-world harvest situations. We decided to perform the analysis on four different size classes, which were extracted from our dataset distribution (Figure 5): small tubers with a volume between 0 and 100 ml, small to medium tubers with a volume between 100 and 150 ml, medium to large tubers with a volume between 150 and 200 ml, and large tubers with a volume between 200 and 500 ml. The ranges were chosen so that a relatively balanced number of tubers ended up in the four classes.
@@
The second practical analysis was on the effect of the potato cultivar on the 3D shape completion result. This analysis provided insights into the degree of generalizability of the 3D completion method at the farm level, as the majority of farmers grow multiple cultivars in a season. We wanted to test if our method can provide accurate volumetric estimates regardless of the potato cultivar. We conducted the experiment with three potato cultivars (Corolle, Kitahime and Sayaka) that differed in shape and size. Corolle was the cultivar with the most elongated and smallest tubers. Kitahime was the cultivar with the most spherical and medium-sized tubers. Sayaka was the cultivar with the largest tubers and the greatest diversity in shape. This may be due to the fact that Sayaka was the cultivar with the most samples in our dataset. As shape metrics, we calculated the elongation factor and the concavity factor. The elongation factor was obtained by dividing the longest dimension of the 3D bounding box of the ground truth mesh by the shortest dimension. Spherical shapes have an elongation factor close to 1 and elongated shapes have an elongation factor closer to 2. The concavity factor was obtained by calculating the Chamfer distance between the original ground truth 3D mesh and the convex hull shape of that mesh. Higher Chamfer distances mean that there are more valleys or concave regions on the surface of the potato tuber.
@@
Our final practical analysis focused on identifying the horizontal image region where ideally the 3D shape completion should be performed. The reason for this analysis is that it is neither computationally desirable nor necessary to analyze all 20 to 30 RGB-D frames of each tuber when it moves over the conveyor belt. The computational efficiency can be improved by performing the 3D shape completion only in the image region with the average lowest RMSE for volumetric estimation. In our analysis, we examined thirteen horizontal image regions (Table 2).
Table 2: Summary of the investigated horizontal image regions.
Region Horizontal image region